OpenStudy (anonymous):

Medal + Fan A researcher wishes to design a fully blocked experiment with groups of subjects representing every possible combination of 1 explanatory variable (control versus treatment) and 4 other variables (male versus female, old versus young, healthy versus ill, and normal weight versus overweight) for a total of 5 variables. A. 32 B. 10 C. 5 D. 16 I know it's either 32 or 16 can someone help?!

OpenStudy (mimi_x3):

5 variables

OpenStudy (mimi_x3):

each have 2 choices each so 2 choices 5 times 2*2*2*2*2=2^5=32